Benny mixed 12 mL of an acidic solution with 11mL of a 41% acidic solution, to make 45% acidic solution. Find the percent acid concentration of the first solution.
Answer by ankor@dixie-net.com(22740)

Benny mixed 12 mL of an acidic solution with 11 mL of a 41% acidic solution, to make 45% acidic solution.
Find the percent acid concentration of the first solution.
:
Let x = the percent of the 12 ml solution in decimal form
12x + .41(11) = .45(12+11)
12x + 4.51 = .45(23)
12x = 10.35 - 4.51
12x = 5.84
x = 5.84/12
x = .4867, 48.67% the percent of the 12 ml solution